Suppose that swine flu (for the purposes of this exercise) affects 1 in 10,000 people in the U.S. The test is known to have a false positive rate of 0.01 -- that is, 1% of all positive tests are actually negative. The probability of a false negative is zero – that is, 100% of all negative tests are truly negative. You test positive. What is the probability that you actually have the swine flu? Hint: Define the event ? as “testing positive” and the event ? as “having the swine flu.”
Answer #1

Let event ? be the event of “testing positive”

Let event ? be the event as “having the swine flu"

Now

P(A) = 0.99 * 1/10,000 + 0.01 * 9999/10,000 = 0.010098

We have to find

P(B/A) = P(A/B) * P(B) / P(A)

0.99* 1/10,000 1 0.0098039 0.010098 102

Thus Probability actually having swine flu = 0.98039% = 0.0098039

Please upvote. Thanks!